# Energy-Enhanced Expansion of the Standard Model Effective Field Theory

### v1

@@ IV Energy Scaling Tables
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}}$ $\lambda^{6}$ $\psi^{2}h^{3}:{1}\;\;$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}}$ $\lambda^{6}$ $h^{2}V^{4}:{1}$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}X^{3}}$ $\lambda^{6}$ $V^{6}:{1}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}X^{4}}$ $\lambda^{8}$ $(\partial V)^{3}V^{2}:{E^{2}}$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{3}}$ $\lambda^{8}$ $(\partial V)^{2}h^{2}V^{2}:{E^{2}},\;\;(\partial V)hV^{4}:{Ev},\;\;V^{6}:{v^{2}}\;\;$
@@
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}H^{2}X^{2}D^{2}}$ $\lambda^{8}$ $h^{2}(\partial^{2}V)V^{3}:{E^{2}},\;\;h(\partial V)V^{4}:{Ev},\;\;V^{6}:{v^{2}}\;\;$
@@
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HX^{2}}$ $\lambda^{8}$ $\psi^{2}h(\partial V)V^{2}:{E^{2}},\;\;\psi^{2}V^{4}:{Ev}\;\;$
${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}\psi^{2}X^{2}D}$ $\lambda^{8}$ $\psi^{2}(\partial V)V^{3}:{E^{2}}$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}X}$ $\lambda^{8}$ $\psi^{2}h^{3}(\partial V):{E^{2}},\;\;\psi^{2}h^{2}V^{2}:{Ev}\;\;$
@@
${\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}\psi^{2}HXD^{2}}$ $\lambda^{8}$ $\psi^{2}h(\partial V)V^{3}:{E^{2}},\;\;\psi^{2}V^{4}:{Ev}\;\;$
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{2}H^{3}D^{2}}$ $\lambda^{8}$ $\psi^{2}h^{3}(\partial V):{E^{2}},\;\;\psi^{2}h^{2}V^{2}:{Ev}\;\;$
@@
${\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}\psi^{4}HD}$ $\lambda^{8}$ $\psi^{4}hV:{E^{2}}\;\;$
@@ V.1 $pp\rightarrow VH(q\bar{q})$
$g_{\bar{q}qV}$ $=\frac{g^{\rm SM}_{\bar{q}qV}}{\lambda^{2}}+\frac{\lambda}{\hat{\Lambda}^{2}}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}+\frac{\lambda^{2}}{\hat{\Lambda}^{2}}\,c_{\psi^{2}H^{2}D}$
$\quad+\frac{\lambda^{5}}{\hat{\Lambda}^{4}}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}H^{3}X}}+\frac{\lambda^{5}}{\hat{\Lambda}^{4}}\,c_{\psi^{2}H^{4}D}+\cdots\,.$ (13)
where we factor out the common dimensionful kinematic factor in order to focus on the scaling. Here, the coefficients $c_{\mathcal{O}}$ (with $\mathcal{O}$ labeling the operator from Tables 3,4) indicate the contribution of the corresponding operator (including the color coding from Sect. III). The factors of $\hat{\Lambda}$ serve as bookkeeping devices to keep track of the order of the various operator contributions, i.e. re-writing $c^{(i)}_{j}$ in Eq. (3) to $c_{j}/{\hat{\Lambda}}^{i}$. In Eq. (V.1), only the first few orders in $\lambda$ are shown.
Figure 1: Topologies for $\bar{q}q\to VH$ production
@@
$g_{hVV}$ $=\frac{g^{\rm SM}_{hVV}}{\lambda}+\frac{\lambda}{\hat{\Lambda}^{2}}\,{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}+\frac{\lambda^{3}}{\hat{\Lambda}^{2}}\,c_{H^{4}D^{2}}$
@@
$g_{\bar{q}qhV}$ $=\frac{\lambda^{2}}{\hat{\Lambda}^{2}}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}+\frac{\lambda^{3}}{\hat{\Lambda}^{2}}\,c_{\psi^{2}H^{2}D}+\frac{\lambda^{4}}{\hat{\Lambda}^{4}}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HXD^{2}}}$
@@
Note that the SM contribution to $g_{hVV}$ carries a weight of $1/\lambda$ since it is proportional to $v$ rather than an energy $E$.
Multiplying the $g_{\bar{q}qV}$ and $g_{hVV}$ expansions and including a factor of $E^{-2}$ which scales $\lambda^{4}$ from the vector boson propagator, the net scaling for the topology on the left in Fig. 1 becomes (to $\mathcal{O}(\lambda^{5})$)
$\mathcal{A}^{(1)}_{qqVh}$ $=\lambda\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{hVV}+\frac{\lambda^{3}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}+\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{hVV}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}$
@@
A comparison of $\mathcal{A}^{(1)}_{qqVh}$ and $\mathcal{A}^{(2)}_{qqVh}$, by counting powers of $\lambda$, indicates that the largest SMEFT effect (i.e. the contribution with the fewest powers of $\lambda$) arises from the right hand topology in Fig 1. However, several refinements are possible. For example, operators whose interference with the SM is suppressed—either because they have a different fermion chirality structure or because they contribute predominantly to transversely polarized vector bosons (in contrast to the predominantly longitudinally polarized $V$ in the SM)—can be neglected. This prescription eliminates the operators shaded in purple, as well as ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}$ and $c_{H^{4}X^{2}}$.
Figure 2: Topologies for $gg\to t\bar{t}H$
Two of the remaining operator classes, $c_{\psi^{2}H^{2}D}$ and ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}$, are tightly constrained by $Z$-pole and Higgs pole measurements Ellis:2018gqa ; Ellis:2020unq ; Bartocci:2023nvp ; Giani:2023gfq . One way to comply with these constraints is to assume that the scale $\Lambda$ is so large that $c\sim\mathcal{O}(1)$ is consistent. This results in a smaller $\lambda$, suppressing all SMEFT effects. However, from the bottom-up perspective, it is also interesting to consider a scenario where $c_{\psi^{2}H^{2}D}$ and ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}$ are small while the other Wilson coefficients remain $\mathcal{O}(1)$. In this case, higher order operators can rise in importance.
@@
$\mathcal{A}^{(1)}_{qqVh}$ $=\lambda\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{hVV}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{hVV}\,g^{\rm SM}_{\bar{q}qV}\,c_{H^{4}D^{2}}+\cdots\,,$
@@ V.2 $gg\to t\bar{t}H$
$\quad+\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{\bar{q}qh}\,{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{X^{3}}}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,g^{\rm SM}_{\bar{q}qh}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}$
@@
$\quad+\frac{\lambda^{6}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,\Bigl{(}g^{\rm SM}_{\bar{q}qh}\,c_{\psi^{2}H^{2}D}+g^{\rm SM}_{\bar{q}qV}\,{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}c_{\psi^{2}H^{3}}}\Bigr{)}+\cdots\,,$ (18)
@@
$\mathcal{A}^{(2)}_{ggttH}$ $=\lambda^{2}\,g^{\rm SM}_{\bar{q}qh}\,\Bigl{(}g^{\rm SM}_{\bar{q}qV}\Bigr{)}^{2}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{\bar{q}qh}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}$
@@
$\quad+\frac{\lambda^{6}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,\Bigl{(}g^{\rm SM}_{\bar{q}qh}\,c_{\psi^{2}H^{2}D}+g^{\rm SM}_{\bar{q}qV}\,{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}c_{\psi^{2}H^{3}}}\Bigr{)}+\cdots\,,$ (19)
$\mathcal{A}^{(3)}_{ggttH}$ $=\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,g^{\rm SM}_{\bar{q}qV}\,{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}+\cdots\,,$ (20)
$\mathcal{A}^{(4)}_{ggttH}$ $=\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,g^{\rm SM}_{\bar{q}qh}\,{\color[rgb]{0.1171875,0.53515625,0.8984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.53515625,0.8984375}c_{\psi^{2}X^{2}D}}+\cdots\,,$ (21)
$\mathcal{A}^{(5)}_{ggttH}$ $=\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,c_{\psi^{2}H^{2}D}+$
$\quad+\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,g^{\rm SM}_{\bar{q}qV}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HXD^{2}}}\cdots\,,$ (22)
$\mathcal{A}^{(6)}_{ggttH}$ $=\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,c_{\psi^{2}H^{2}D}$
@@
$\quad+\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,\Bigl{(}g^{\rm SM}_{VVV}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HXD^{2}}}+{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{X^{3}}}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}\Bigr{)}+\cdots\,,$ (23)
$\mathcal{A}^{(7)}_{ggttH}$ $=\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}+\cdots\,,$ (24)
$\mathcal{A}^{(8)}_{ggttH}$ $=\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX}}+\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,{\color[rgb]{0.55859375,0.140625,0.66796875}\definecolor[named]{pgfstrokecolor}{rgb}{0.55859375,0.140625,0.66796875}c_{\psi^{2}HX^{2}}}\,.$ (25)
@@
As in the $pp\to VH$ example, one can explore how different assumptions affect the relative importance of the SMEFT terms. For instance, if one drops the suppressed (or non-interfering) terms, the lowest-order SMEFT contribution originates from the four-point vertex proportional to $\lambda^{5}\,{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}$ in the topologies labeled a.) and c.) in Fig. 2. However, since operators of the ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}$ class are constrained by Higgs-pole measurements, setting ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{H^{2}X^{2}}}$ to zero shifts the leading SMEFT contributions to $\mathcal{O}(\lambda^{6})$, five-particle interaction which arises from the dimension-eight operator ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{\psi^{2}X^{2}D}}$.
@@
Alternatively, if we zero all operators generated at loop level, then the leading SMEFT terms are $\lambda^{5}\,c_{\psi^{2}H^{2}D}$ and $\lambda^{6}\,{\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{\psi^{2}H^{3}}}$. If these are somehow suppressed – given that $c_{\psi^{2}H^{2}D}$ and ${\color[rgb]{0.90234375,0.31640625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.31640625,0}c_{\psi^{2}H^{3}}}$ are bounded by $Z$-pole and Higgs pole measurements, respectively – the next terms in the $\lambda$ expansion (all $\mathcal{O}(\lambda^{7})$) are $c_{H^{4}D^{2}}$ at dimension six and $c_{\psi^{2}H^{2}D^{3}},c_{\psi^{2}H^{2}XD},c_{\psi^{2}H^{4}D}$ at dimension eight.

### v2

@@ IV Energy Scaling Tables
$\underline{\vphantom{y}\psi^{2}H^{3}}$ $\lambda^{6}$ $\psi^{2}h^{3}:{1}\;\;$
$\bm{H^{2}X^{2}}$ $\lambda^{6}$ $h^{2}V^{4}:{1}$
$\bm{X^{3}}$ $\lambda^{6}$ $V^{6}:{1}$
@@
$\bm{X^{4}}$ $\lambda^{8}$ $(\partial V)^{2}V^{4}:{E^{2}}$
@@
$\bm{H^{2}X^{3}}$ $\lambda^{8}$ $(\partial V)^{2}h^{2}V^{2}:{E^{2}},\;\;(\partial V)hV^{4}:{Ev},\;\;V^{6}:{v^{2}}\;\;$
@@
$\bm{H^{2}X^{2}D^{2}}$ $\lambda^{8}$ $h^{2}(\partial^{2}V)V^{3}:{E^{2}},\;\;h(\partial V)V^{4}:{Ev},\;\;V^{6}:{v^{2}}\;\;$
@@
$\underline{\vphantom{y}\bm{\psi^{2}HX^{2}}}$ $\lambda^{8}$ $\psi^{2}h(\partial V)V^{2}:{E^{2}},\;\;\psi^{2}V^{4}:{Ev}\;\;$
$\bm{\psi^{2}X^{2}D}$ $\lambda^{8}$ $\psi^{2}(\partial V)V^{3}:{E^{2}}$
$\underline{\vphantom{y}\psi^{2}H^{3}X}$ $\lambda^{8}$ $\psi^{2}h^{3}(\partial V):{E^{2}},\;\;\psi^{2}h^{2}V^{2}:{Ev}\;\;$
@@
$\underline{\vphantom{y}\bm{\psi^{2}HXD^{2}}}$ $\lambda^{8}$ $\psi^{2}h(\partial V)V^{3}:{E^{2}},\;\;\psi^{2}V^{4}:{Ev}\;\;$
$\underline{\vphantom{y}\psi^{2}H^{3}D^{2}}$ $\lambda^{8}$ $\psi^{2}h^{3}(\partial V):{E^{2}},\;\;\psi^{2}h^{2}V^{2}:{Ev}\;\;$
@@
$\underline{\vphantom{y}\psi^{4}HD}$ $\lambda^{8}$ $\psi^{4}hV:{E^{2}}\;\;$
@@ V.1 $pp\rightarrow VH(q\bar{q})$
$g_{\bar{q}qV}$ $=\frac{g^{\rm SM}_{\bar{q}qV}}{\lambda^{2}}+\frac{\lambda}{\hat{\Lambda}^{2}}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}+\frac{\lambda^{2}}{\hat{\Lambda}^{2}}\,c_{\psi^{2}H^{2}D}$
$\quad+\frac{\lambda^{5}}{\hat{\Lambda}^{4}}\,\underline{\vphantom{y}\bm{c_{\psi^{2}H^{3}X}}}+\frac{\lambda^{5}}{\hat{\Lambda}^{4}}\,c_{\psi^{2}H^{4}D}+\cdots\,.$ (13)
where we factor out the common dimensionful kinematic factor in order to focus on the scaling. Here the factor $1/\lambda^{2}$ originates from the two external on-shell spinor’s wave-function normalization, each supplying $\sqrt{E}\sim\lambda^{-1}$ in our power counting. The coefficients $c_{\mathcal{O}}$ (with $\mathcal{O}$ labeling the operator from Tables 3,4) indicate the contribution of the corresponding operator (including the coding from Sect. III). The factors of $\hat{\Lambda}$ serve as bookkeeping devices to keep track of the order of the various operator contributions, i.e. re-writing $c^{(i)}_{j}$ in Eq. (3) to $c_{j}/{\hat{\Lambda}}^{i}$. In Eq. (V.1), only the first few orders in $\lambda$ are shown.
[…]
@@
$g_{hVV}$ $=\frac{g^{\rm SM}_{hVV}}{\lambda}+\frac{\lambda}{\hat{\Lambda}^{2}}\,\bm{c_{H^{2}X^{2}}}+\frac{\lambda^{3}}{\hat{\Lambda}^{2}}\,c_{H^{4}D^{2}}$
@@
$g_{\bar{q}qhV}$ $=\frac{\lambda^{2}}{\hat{\Lambda}^{2}}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}+\frac{\lambda^{3}}{\hat{\Lambda}^{2}}\,c_{\psi^{2}H^{2}D}+\frac{\lambda^{4}}{\hat{\Lambda}^{4}}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HXD^{2}}}}$
@@
Note that the SM contribution to $g_{hVV}$ carries a weight of $1/\lambda$ since it is proportional to $v$ rather than an energy $E$. Moreover, we include sub-leading vertices, such as those that scale as $\lambda^{3}$ and $\lambda^{5}$ since in this case they end up providing the leading contributions once combined in the amplitude. These sub-leading vertices are not presented in the tables but can be simply determined from the counting algorithm.
Multiplying the $g_{\bar{q}qV}$ and $g_{hVV}$ expansions and including a factor of $E^{-2}$ which scales as $\lambda^{4}$ from the vector boson propagator, the net scaling for the topology on the left in Fig. 1 becomes (to $\mathcal{O}(\lambda^{5})$)
$\mathcal{A}^{(1)}_{qqVh}$ $=\lambda\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{hVV}+\frac{\lambda^{3}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,\bm{c_{H^{2}X^{2}}}+\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{hVV}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}$
@@
A comparison of $\mathcal{A}^{(1)}_{qqVh}$ and $\mathcal{A}^{(2)}_{qqVh}$, by counting powers of $\lambda$, indicates that the largest SMEFT effect (i.e. the contribution with the fewest powers of $\lambda$) arises from the right hand topology in Fig 1. However, several refinements are possible. For example, operators whose interference with the SM is suppressed—either because they have a different fermion chirality structure or because they contribute predominantly to transversely polarized vector bosons (in contrast to the predominantly longitudinally polarized $V$ in the SM)—can be neglected. This prescription eliminates the operators shaded in purple, as well as $\bm{c_{H^{2}X^{2}}}$ and $c_{H^{4}X^{2}}$.
[…]
Two of the remaining operator classes, $c_{\psi^{2}H^{2}D}$ and $\bm{c_{H^{2}X^{2}}}$, are tightly constrained by $Z$-pole and Higgs pole measurements Ellis:2018gqa ; Ellis:2020unq ; Bartocci:2023nvp ; Giani:2023gfq . One way to comply with these constraints is to assume that the scale $\Lambda$ is so large that $c\sim\mathcal{O}(1)$ is consistent. This results in a smaller $\lambda$, suppressing all SMEFT effects. However, from the bottom-up perspective, it is also interesting to consider a scenario where $c_{\psi^{2}H^{2}D}$ and $\bm{c_{H^{2}X^{2}}}$ are small while the other Wilson coefficients remain $\mathcal{O}(1)$. In this case, higher order operators can rise in importance.
@@
$\mathcal{A}^{(1)}_{qqVh}$ $=\lambda\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{hVV}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,c_{H^{4}D^{2}}+\cdots\,,$
@@ V.2 $gg\to t\bar{t}H$
$\quad+\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{\bar{q}qh}\,\bm{c_{X^{3}}}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,g^{\rm SM}_{\bar{q}qh}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}$
@@
$\quad+\frac{\lambda^{6}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,\Bigl{(}g^{\rm SM}_{\bar{q}qh}\,c_{\psi^{2}H^{2}D}+g^{\rm SM}_{\bar{q}qV}\,\underline{\vphantom{y}c_{\psi^{2}H^{3}}}\Bigr{)}+\cdots\,,$ (18)
@@
$\mathcal{A}^{(2)}_{ggttH}$ $=\lambda^{2}\,g^{\rm SM}_{\bar{q}qh}\,\Bigl{(}g^{\rm SM}_{\bar{q}qV}\Bigr{)}^{2}+\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,g^{\rm SM}_{\bar{q}qh}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}$
@@
$\quad+\frac{\lambda^{6}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,\Bigl{(}g^{\rm SM}_{\bar{q}qh}\,c_{\psi^{2}H^{2}D}+g^{\rm SM}_{\bar{q}qV}\,\underline{\vphantom{y}c_{\psi^{2}H^{3}}}\Bigr{)}+\cdots\,,$ (19)
$\mathcal{A}^{(3)}_{ggttH}$ $=\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,g^{\rm SM}_{\bar{q}qV}\,\bm{c_{H^{2}X^{2}}}+\cdots\,,$ (20)
$\mathcal{A}^{(4)}_{ggttH}$ $=\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,g^{\rm SM}_{\bar{q}qh}\,\underline{\vphantom{y}c_{\psi^{2}X^{2}D}}+\cdots\,,$ (21)
$\mathcal{A}^{(5)}_{ggttH}$ $=\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}+\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,g^{\rm SM}_{\bar{q}qV}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HXD^{2}}}}+\cdots\,,$ (22)
$\mathcal{A}^{(6)}_{ggttH}$ $=\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{VVV}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}$
@@
$\quad+\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,\Bigl{(}g^{\rm SM}_{VVV}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HXD^{2}}}}+\bm{c_{X^{3}}}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}\Bigr{)}+\cdots\,,$ (23)
$\mathcal{A}^{(7)}_{ggttH}$ $=\frac{\lambda^{5}}{\hat{\Lambda}^{2}}\,g^{\rm SM}_{\bar{q}qV}\,\bm{c_{H^{2}X^{2}}}+\cdots\,,$ (24)
$\mathcal{A}^{(8)}_{ggttH}$ $=\frac{\lambda^{4}}{\hat{\Lambda}^{2}}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX}}}+\frac{\lambda^{6}}{\hat{\Lambda}^{4}}\,\underline{\vphantom{y}\bm{c_{\psi^{2}HX^{2}}}}\,.$ (25)
@@
As in the $pp\to VH$ example, one can explore how different assumptions affect the relative importance of the SMEFT terms. For instance, if one drops the suppressed (or non-interfering) terms, the lowest-order SMEFT contribution originates from the four-point vertex proportional to $\lambda^{5}\,\bm{c_{H^{2}X^{2}}}$ in the topologies labeled a.) and c.) in Fig. 2. However, since operators of the $\bm{c_{H^{2}X^{2}}}$ class are constrained by Higgs-pole measurements, setting $\bm{c_{H^{2}X^{2}}}$ to zero shifts the leading SMEFT contributions to $\mathcal{O}(\lambda^{6})$, five-particle interaction which arises from the dimension-eight operator $\bm{c_{\psi^{2}X^{2}D}}$.
@@
Alternatively, if we zero all operators generated at loop level, then the leading SMEFT terms are $\lambda^{5}\,c_{\psi^{2}H^{2}D}$ and $\lambda^{6}\,\bm{c_{\psi^{2}H^{3}}}$. If these are somehow suppressed – given that $c_{\psi^{2}H^{2}D}$ and $\bm{c_{\psi^{2}H^{3}}}$ are bounded by $Z$-pole and Higgs pole measurements, respectively – the next terms in the $\lambda$ expansion (all $\mathcal{O}(\lambda^{7})$) are $c_{H^{4}D^{2}}$ at dimension six and $c_{\psi^{2}H^{2}D^{3}},c_{\psi^{2}H^{2}XD},c_{\psi^{2}H^{4}D}$ at dimension eight.